.01 moles of gaseous HCl is added to 100-ml of the buffer composed of .5M sodium acetate and .5M acetic acid. Find the pH
Answer #1

Apply buffer equation

pH = pKa +loc(acetate/acid)

pH = 4.75 +log(0.5/0.5) = 4.75

after adding

0.01 mol = 100 ml = 0.01*100 = 1 mmol of H+

then

mmol of conjugate reacted = 0.5*100 -1 = 50-1 = 49

mmo lof acid formed = 0.5*100 + 1 = 51

then

pH = pKa +log(acetate/acid)

pH = 4.75+log(49/51) = 4.732625